OpenStudy (firechild17):

A pharmacy claims that the average medication costs $32 but it could differ as much as $8. Write and solve an absolute value inequality to determine the range of medication costs at this pharmacy. |x − 32| ≥ 8; The medication costs range from $24 to $40 |x − 32| ≥ 8; The medications cost less than $24 or greater than $40. |x − 32| ≤ 8; The medication costs range from $24 to $40 |x − 32| ≤8; The medications cost less than $24 or greater than $40.

OpenStudy (firechild17):

@563blackghost

563blackghost (563blackghost):

First identify the key words....we see that it says `costs $32 but it could differ as much as $8.`The keyword here is `as much as` which would be \(\large{\le}\).

OpenStudy (firechild17):

ok

OpenStudy (firechild17):

so thats greater than or equal to right?

563blackghost (563blackghost):

well it would actually be `less than or equal to`.

OpenStudy (firechild17):

oh i get those confused

563blackghost (563blackghost):

Now to find x we would split this into two equation with the outcome a negative and a positive... \(\huge\bf{x-32 \le 8}\) \(\huge\bf{x -32 \le -8}\) Simplify the two.

563blackghost (563blackghost):

Im sorry but I have to go :( maybe @3mar or @StudyGurl14 can help :)

563blackghost (563blackghost):

The process is suppose to be typed as... \(\huge\bf{-8 \le x-32 \le 8}\) Sorry I typed it wrong

OpenStudy (firechild17):

\[X \le 40\]

OpenStudy (firechild17):

how do i solve that?

OpenStudy (firechild17):

@3mar this is the question but i have to go for dinner Can we finish this tomorrow?

OpenStudy (3mar):

May be In Sha' Allah Just let me know tomorrow audi3omar@Skype audi3mar@Yahoo

OpenStudy (firechild17):

ok

OpenStudy (firechild17):

@3mar

OpenStudy (firechild17):

im back from dinner

OpenStudy (firechild17):

@3mar this one

OpenStudy (3mar):

where?

OpenStudy (firechild17):

its at the top...

OpenStudy (3mar):

Where are you stuck?

OpenStudy (firechild17):

I think its B but im unsure

OpenStudy (3mar):

Ok

OpenStudy (firechild17):

am i right?

OpenStudy (3mar):

Let me check

OpenStudy (firechild17):

k

OpenStudy (3mar):

It would be D. This is because the difference in the cost doesn't exceed 8$, which means the cost difference is less or equal to 8

OpenStudy (3mar):

Did you get it?

OpenStudy (3mar):

my 10 min is up!

563blackghost (563blackghost):

Actually it would be C.....You would simplify the expression by adding 32 to each side.... \(\huge\bf{24 \le x \le 40}\) `The medication costs range from $24 to $40`

OpenStudy (firechild17):

ok sorry

563blackghost (563blackghost):

No its alright its best to know :)

OpenStudy (firechild17):

i need help with bio